# MLLM-CL: Continual Learning for Multimodal Large Language Models

### v1

@@ Abstract
Recent Multimodal Large Language Models (MLLMs) excel in vision-language understanding but face challenges in adapting to dynamic real-world scenarios that require continuous integration of new knowledge and skills. While continual learning (CL) offers a potential solution, existing benchmarks and methods suffer from critical limitations. In this paper, we introduce MLLM-CL, a novel benchmark encompassing domain and ability continual learning, where the former focuses on independently and identically distributed (IID) evaluation across evolving mainstream domains, whereas the latter evaluates on non-IID scenarios with emerging model ability. Methodologically, we propose preventing catastrophic interference through parameter isolation, along with an MLLM-based routing mechanism. Extensive experiments demonstrate that our approach can integrate domain-specific knowledge and functional abilities with minimal forgetting, significantly outperforming existing methods.
@@ 1 Introduction
Recent advancements in Multimodal Large Language Models (MLLMs) liu2023improvedllava ; chen2024allava have demonstrated remarkable capabilities in vision-language understanding. These models typically undergo supervised finetuning on carefully curated multi-task datasets, whereas real-world applications require continuous adaptation to evolving user requirements and dynamic data streams with shifting domain distributions. To incorporate new knowledge and skills, full retraining of large models is costly in both time and computing resources, besides straightforward finetuning on novel tasks often results in catastrophic forgetting mccloskey1989catastrophic ; zhai2023investigating . Therefore, for deployment in ever-changing environments, there is an urgent need to develop MLLMs capable of continually consolidating new skills while maintaining performance on prior tasks. Recently, a few studies chen2024coin ; zeng2024modalprompt ; cao2024continual ; guo2025hide ; he2023continual have explored continual learning (CL) of MLLMs. However, current works still have key limitations in both benchmarks and methodologies, preventing them from effectively exploring CL in MLLMs.
@@
Firstly, there is a lack of well-established benchmarks. Chen et al., chen2024coin proposed the first continual instruction tuning benchmark for MLLMs comprising several downstream datasets, while some of them have already been learned during the early supervised finetuning (SFT) phase of MLLM. Huai et al., huai2025cl divided VQAv2 goyal2017making into several tasks and conducted continual instruction tuning directly from the LLaVA liu2023llava base model. However, in real-world applications, continually learning subsets of a specific dataset is impractical and it is unlikely to finetune an MLLM on downstream tasks without any SFT on general multimodal data. Moreover, those benchmarks only consider independently and identically distributed (IID) evaluation (the training and test sets are split from the same dataset), while the model would encounter non-IID inputs in practice.
@@
Secondly, existing methods have notable limitations: (1) Some approaches share the same set of parameters for different tasks chen2024coin ; huang2024class . This might be suitable for a conventional class-incremental learning scenario where different tasks often belong to the same dataset. However, MLLMs often encounter inputs from various domains, and the inherent task conflicts wei2025asymlora ; yang2024low would lead to loss of plasticity during continual learning, particularly when handling heterogeneous modalities across divergent domains. (2) Parameter isolation methods have to determine which task-specific parameters to apply for a given input during inference. This selection is usually driven by simple hand-crafted similarity metrics zeng2024modalprompt ; guo2025hide , which can be unreliable when confronted with complex multimodal data, consequently undermining overall performance.
@@
In this paper, we establish a novel benchmark MLLM-CL, which includes two practical settings, i.e., domain continual learning (DCL) and ability continual learning (ACL), as shown in Fig. 1. Specifically, DCL aims to equip the model with domain-specific knowledge continually by learning and evaluating on several mainstream domains (remote sensing, medical, autonomous driving, science and finance), where the training and test sets are IID. Differently, ACL focuses on incorporating fundamental abilities (OCR, math & logic, visual perception and GUI agent), which are evaluated on non-IID test sets. Together, these two settings provide a comprehensive and realistic evaluation for continual learning of MLLMs.
@@
Further, we design a novel method called MLLM-enhanced Router with Low-Rank Adaptation (MR-LoRA) to build an efficient, lifelong-evolving MLLM. For plasticity preservation, we employ domain or ability-specific LoRA modules hu2021lora that maintain parameter isolation across sequentially arriving tasks, enabling comprehensive acquisition of new knowledge while preventing catastrophic interference through explicit architectural decoupling. Concurrently, to enhance parameter selection accuracy in complex multimodal scenarios, we devise a multimodal routing mechanism that leverages the model’s intrinsic multimodal understanding capabilities to automatically align input patterns with optimal task parameters. This strategy effectively transforms the MLLM’s knowledge into an explicit expert selector.
@@ 2 Related Work
2.1 Continual Learning
Continual learning addresses the challenge of training machine learning models to acquire new knowledge incrementally while retaining performance on previously learned tasks, a capability critical for real-world applications where data distributions evolve over time kirkpatrick2017overcoming . Researchers have primarily developed four main strategies for continual learning: rehearsal-based methods lavda2018continual ; zhu2021prototype ; buzzega2020dark , regularization-based methods kirkpatrick2017overcoming ; li2017learning ; aljundi2018memory , structure-based methods mallya2018piggyback ; Douillard_Rame_Couairon_Cord_2022 , and prompt-based methods l2p ; smith2023coda . These strategies can be combined to improve performance further.
Continual learning in large language models has recently gained much attention wu2024continual ; shi2024continual . According to the training stages, we can divide them into continual pre-training jang2022temporalwiki ; jang2021towards ; cossu2024continual , continual instruction tuning razdaibiedina2023progressive ; zan2022cert ; yin2022contintin ; wang2023orthogonal and continual alignment zhang2024cppo ; suhr2024continual . However, few studies focus on continual learning of MLLMs chen2024coin ; zeng2024modalprompt ; cao2024continual ; guo2025hide . These prior attempts establish benchmarks with a simple dataset incremental setting where the training and test sets are independently and identically distributed. There are some works that focus on conducting continual instruction tuning directly from the model after the pre-training process huai2025cl ; he2023continual . While these efforts have advanced the development of continual learning for MLLMs to some extent, they exhibit an apparent gap with the real-world production environment. Therefore, our work fills this gap and proposes a comprehensive and practical benchmark including adding domain specific knowledge (IID testing) and general abilities (non-IID testing) for continual learning of multimodal large language models.
2.2 Multimodal Large Language Models
Recent advances in MLLMs have demonstrated remarkable capabilities in multimodal understanding, open-ended generation, and instruction following across modalities. Early efforts, such as BLIP li2022blip ; li2023blip , LLaVA liu2023llava ; liu2023improvedllava and Qwen-VL bai2023qwen use image encoders radford2021learning and projectors to transfer multimodal inputs into language embedding space. LLaVA-NeXT liu2024llavanext , InternVL chen2024internvl and Cambrian1 tong2024cambrian utilize richer image resolution techniques and more diverse SFT data to achieve better performance. Recent advances, such as GPT4o gpt4o , LLaVA-OneVision li2024llava , Qwen2.5-VL Qwen2.5-VL , VITA fu2024vita ; fu2025vita expand the ability of MLLM into more modalities such as video, audio, etc. With the rapid growth of MLLMs, the costs associated with training from scratch have increased dramatically li2024llava ; tong2024cambrian ; Qwen2.5-VL ; chen2024expanding . Therefore, adapting MLLMs to dynamic environments by retraining them from scratch becomes expensive and inefficient, creating an imperative demand for continual learning of MLLMs.
@@ 3 MLLM-CL Benchmark
In this section, we provide the problem formulation and introduce the continual learning benchmark MLLM-CL. Based on the general ability and domain-specific knowledge updated in the instruction tuning stage, we divide our benchmark into domain continual learning and ability continual learning, respectively. In domain continual learning, we desire the model to learn knowledge continually and the training sets and the test sets are IID. While in ability continual learning, we desire the model to enhance different abilities from the training data and generalize to non-IID test sets.
@@
Table 1: Statistics of the training datasets and test datasets for domain continual learning and ability continual learning. In domain continual learning, "RS" stands for remote sensing, "Med" is medical, "AD" is autonomous driving, "Sci" stands for science, "Fin" means finance. In ability continual learning, "M & L" stands for math & logic. "VP" means visual perception.
@@
Figure 2: Data pipeline for generating StockQA instruction tuning dataset.
@@
Domain Continual Learning (DCL). Continually adding domain knowledge is crucial for constructing a powerful MLLM. To achieve this goal, we propose domain continual learning and choose five mainstream and common domains: remote sensing, medical, science, autonomous driving and finance. Specifically, we choose RSVQA lobry2020rsvqa , PathVQA he2020pathvqa , DriveLM sima2023drivelm , FinVis wang2023finvis , AI2D kembhavi2016diagram , SciVerse guo2025sciverse , MapQA chang2022mapqa and TQA kembhavi2017you . However, we find FinVis is a caption dataset and the description is generated by GPT4 achiam2023gpt in Chinese, which may result to a language gap and is not convenient for evaluation. Therefore, we regenerate the SFT and test data as multi-choice questions (MCQ) and yes-or-no questions (Y/N) using a questioner-inspector data pipeline. Fig. 2 shows the overall data pipeline. We use two agents, a QA generator and an inspector. Considering the varying task difficulties, we use Qwen2.5-VL-72b Qwen2.5-VL to generate multiple choice QA pairs and Qwen2.5-VL-7b Qwen2.5-VL to generate Y/N QA pairs. For the inspector, we use Qwen2.5-VL-7b Qwen2.5-VL to check the correctness of each QA pair. After checking by the inspector, we use a rule-based formatting to get the final dataset and name it StockQA. All experiments are conducted using the vllm kwon2023efficient engine. Please refer to the supplementary materials to see the detailed prompts for each agent, the rules for filtering, examples and statistics of StockQA dataset. Tab. 1 shows the statistics of the datasets for domain continual learning in the MLLM-CL benchmark and Fig. 1 shows some examples of DCL. More examples are provided in the supplementary materials.
@@
Ability Continual Learning (ACL). Domain continual learning assumes that training and test data are IID. However, achieving IID between training and test sets is often challenging in real-world scenarios, which has been ignored by existing benchmarks chen2024coin ; zeng2024modalprompt ; guo2025hide ; cao2024continual . Therefore, we consider a more challenging setting with non-IID training and test data, which we term ability continual learning. For ability continual learning, we select four fundamental abilities for the MLLM to learn sequentially: OCR, math & logic, visual perception and GUI agent. In terms of the SFT data, we collect the relevant data from LLaVA-OneVision li2024llava , Monkey li2023monkey , ScreenQA hsiao2022screenqa , Screen2Words wang2021screen2words , MultiUI liu2024harnessing , Math-LLaVA shihu2024mathllava , MAVIS zhang2024mavismathematicalvisualinstruction , CLVER johnson2017clevr and TallyQA acharya2019tallyqa and test them on OCRBench OCRBench , MathVista lu2024mathvista , MMTBench-GUI mmtbench and CV-Bench-Counting tong2024cambrian , respectively. Tab. 1 presents the details of the datasets for training and test in ability continual learning and Fig. 1 illustrates a demonstration of ACL. Additional examples can be found in the supplementary materials.
@@ 4 The Proposed Method: MR-LoRA
4.1 Expert Learning without Task Conflict
Figure 3: Comparison of new task performance: (a) Average new task accuracy on five domains. (b) Performance on Sci after learning AD. (d) Extending training cannot fully solve the loss of plasticity.
[…]
Learning Low-Rank Expert. To circumvent inter-task conflicts when learning new domains, we propose initializing a fresh LoRA module from scratch for each task. Compared to the original parameters of the large model, LoRA introduces minimal additional parameters, enabling domain-specific adaptation via lightweight, task-exclusive adapters.
@@ 4.1 Expert Learning without Task Conflict
4.2 Router Selection with MLLM
Few-shot Router Tuning. In our framework, we tune a low-rank expert for each domain or capability, and dynamically select the most appropriate expert at inference time. While existing selection strategies zeng2024modalprompt ; guo2025hide rely on simple similarity measures, e.g., computing cosine similarity between task prototypes and sample features in the embedding space, multimodal scenarios involve more complex inputs. Therefore, we propose leveraging the MLLM’s intrinsic capability to process complex multimodal inputs by tuning an MLLM-based selection router. This router identifies the corresponding expert for each input. Specifically, for each task, we collect a few-shot set $\mathcal{M}_{t}=\{(x^{\text{img}}_{t,i},x^{\text{ins}}_{t,i})\}_{i=1}^{m}$, where $m\ll N_{t}$ (we set $m=20$ in all experiments). After each continual learning phase, the accumulated few-shot data $\{\mathcal{M}_{j}\}_{j=1}^{t}$ and expert model descriptions are transformed into structured instructions. We adopt a generative style to select the most suitable expert and tune the MLLM using a router LoRA via autoregressive loss liu2023improvedllava . An illustration of the router selection prompt for domain continual learning is provided in Fig. 4.
The proposed MLLM-based routing mechanism offers notable advantages: (1) The MLLM’s strong multimodal understanding capacity ensures robust expert selection performance on complex multimodal inputs. (2) The selection router is parameter-efficient and learned with few-shot unlabeled image-question pairs, allowing on-the-fly adaptation.

### v2

@@ Abstract
Recent Multimodal Large Language Models (MLLMs) excel in vision-language understanding but face challenges in adapting to dynamic real-world scenarios that require continuous integration of new knowledge and skills. While continual learning (CL) offers a potential solution, existing benchmarks and methods suffer from critical limitations. In this paper, we introduce MLLM-CL, a novel benchmark encompassing domain and ability continual learning, where the former focuses on independently and identically distributed (IID) evaluation across evolving mainstream domains, whereas the latter evaluates on non-IID scenarios with new model abilities. Methodologically, we propose preventing catastrophic interference through parameter isolation and an MLLM-based routing mechanism. Extensive experiments demonstrate that our approach can integrate domain-specific knowledge and functional abilities with minimal forgetting, significantly outperforming existing methods. Our benchmark and code are available at https://github.com/bjzhb666/MLLM-CL.
@@ 1 Introduction
Recent advancements in Multimodal Large Language Models (MLLMs) (Liu et al., 2024a; Chen et al., 2024b) have demonstrated remarkable capabilities in vision-language understanding. These models typically undergo supervised finetuning on carefully curated multi-task datasets, whereas real-world applications require continuous adaptation to evolving user requirements and dynamic data streams with shifting domain distributions. To incorporate new knowledge and skills, full retraining of large models is costly in both time and computing resources; besides, straightforward finetuning on novel tasks often results in catastrophic forgetting (McCloskey & Cohen, 1989; Zhai et al., 2023). Therefore, for deployment in ever-changing environments, there is an urgent need to develop MLLMs capable of continually consolidating new skills while maintaining performance on prior tasks. Recently, a few studies (Chen et al., 2024a; Zeng et al., 2024; Cao et al., 2024; Guo et al., 2025a; He et al., 2023) have explored continual learning (CL) of MLLMs. However, current works still have key limitations in both benchmarks and methodologies, preventing them from effectively exploring CL in MLLMs.
@@
Firstly, there is a lack of well-established benchmarks. Chen et al. (2024a) proposed the first continual instruction tuning benchmark for MLLMs comprising several downstream datasets, while some of them have already been learned during the early supervised finetuning (SFT) phase of MLLM. Huai et al. (2025) divided VQAv2 (Goyal et al., 2017) into several tasks and conducted continual instruction tuning directly from the LLaVA (Liu et al., 2023) base model. However, in real-world applications, continually learning subsets of a specific dataset is impractical, and it is unlikely to finetune an MLLM on downstream tasks without any SFT on general multimodal data. Moreover, those benchmarks only consider independently and identically distributed (IID) evaluation (the training and test sets are split from the same dataset), while the model would encounter non-IID inputs in practice.
@@
Secondly, existing methods have notable limitations: (1) Some approaches share the same set of parameters for different tasks (Chen et al., 2024a; Huang et al., 2024). This might be suitable for a conventional class-incremental learning scenario where different tasks often belong to the same dataset. However, MLLMs often encounter inputs from various domains, and the inherent task conflicts (Wei et al., 2025; Yang et al., 2024) would lead to loss of plasticity during continual learning, particularly when handling heterogeneous modalities across divergent domains. (2) Parameter isolation methods have to determine which task-specific parameters to apply for a given input during inference. This selection is usually driven by simple hand-crafted similarity metrics (Zeng et al., 2024; Guo et al., 2025a), which can be unreliable when confronted with complex multimodal data, consequently undermining overall performance.
@@
In this paper, we establish a novel benchmark MLLM-CL, which includes two practical settings, i.e., domain continual learning (DCL) and ability continual learning (ACL), as shown in Fig. 1. Specifically, DCL aims to equip the model with domain-specific knowledge continually by learning and evaluating on several mainstream domains (remote sensing, medical, autonomous driving, science, and finance), where the training and test sets are IID. Differently, ACL focuses on incorporating fundamental abilities (OCR, math & logic, visual perception, and GUI agent), which are evaluated on non-IID test sets. Together, these two settings provide a comprehensive and realistic evaluation for continual learning of MLLMs.
@@
Further, we design a novel method to build an efficient, lifelong-evolving MLLM. For plasticity preservation, we employ domain or ability-specific Low-Rank Adaptation (LoRA) modules (Hu et al., 2021) that maintain parameter isolation across sequentially arriving tasks, enabling comprehensive acquisition of new knowledge while preventing catastrophic interference through explicit architectural decoupling. Concurrently, to enhance parameter selection accuracy in complex multimodal scenarios, we devise a multimodal routing mechanism that leverages the model’s intrinsic multimodal understanding capabilities to automatically align input patterns with optimal task parameters. This strategy effectively transforms the MLLM’s knowledge into an explicit expert selector.
@@ 2 Related Work
Continual Learning.
Researchers have developed primarily four main strategies for continual learning: rehearsal-based methods (Lavda et al., 2018; Buzzega et al., 2020), regularization-based methods (Kirkpatrick et al., 2017; Li & Hoiem, 2017), structure-based methods (Mallya et al., 2018; Douillard et al., 2022), and prompt-based methods (Wang et al., 2022; Smith et al., 2023). CL in large language models has recently gained much attention (Wu et al., 2024; Shi et al., 2024a). According to the training stages, we can divide them into continual pre-training (Jang et al., 2022; Cossu et al., 2024), continual instruction tuning (Razdaibiedina et al., 2023; Zan et al., 2022; Yin et al., 2022; Wang et al., 2023a), and continual alignment (Zhang et al., 2024a; Suhr & Artzi, 2024). However, few studies focus on continual learning of MLLMs (Chen et al., 2024a; Zeng et al., 2024; Cao et al., 2024; Guo et al., 2025a; c). These prior attempts establish benchmarks with a simple dataset incremental setting where training and test sets are distributed independently and identically. Some works focus on conducting continuous instruction tuning directly from the model after the pretraining process (Huai et al., 2025; He et al., 2023). While these efforts have advanced the development of continual learning for MLLMs to some extent, they exhibit an apparent gap with the real-world production environment. Therefore, our work fills this gap and proposes a comprehensive and practical benchmark, including adding domain-specific knowledge and general abilities for CL of MLLM.
Multimodal Large Language Models.
Recent advances in MLLMs have demonstrated remarkable capabilities in multimodal understanding, open-ended generation, and instruction following across modalities. Early efforts, such as LLaVA (Liu et al., 2023; 2024a) and Qwen-VL (Bai et al., 2023), use image encoders (Radford et al., 2021) and projectors to transfer multimodal inputs into language embedding space. Recent advances (OpenAI, 2024; Li et al., 2024a; Bai et al., 2025; Fu et al., 2025) expand the ability of MLLM into more modalities, such as video and audio. With the rapid growth of MLLMs, the costs associated with training from scratch have increased dramatically (Li et al., 2024a; Tong et al., 2024; Bai et al., 2025; Chen et al., 2024c). Therefore, adapting MLLMs to dynamic environments by retraining them from scratch becomes expensive and inefficient, creating an imperative demand for continual learning of MLLMs.
@@ 3 MLLM-CL Benchmark
In this section, we provide the problem formulation and introduce the continual learning benchmark MLLM-CL. Based on the general ability and domain-specific knowledge updated in the instruction tuning stage, we divide our benchmark into domain continual learning and ability continual learning, respectively. In domain continual learning, we desire the model to learn knowledge continually, and the training sets and the test sets are IID. While in ability continual learning, we desire the model to enhance different abilities from the training data and generalize to non-IID test sets.
@@
Table 1: Statistics of the training datasets and test datasets for domain continual learning and ability continual learning. In domain continual learning, "RS" stands for remote sensing, "Med" is medical, "AD" is autonomous driving, "Sci" stands for science, and "Fin" means finance. In ability continual learning, "M & L" stands for math & logic. "VP" means visual perception.
@@
Domain Continual Learning (DCL). Continually adding domain knowledge is crucial for constructing a powerful MLLM. To achieve this goal, we propose domain continual learning and choose five mainstream and common domains: remote sensing, medical, science, autonomous driving, and finance. Specifically, we choose RSVQA (Lobry et al., 2020), PathVQA (He et al., 2020), DriveLM (Sima et al., 2023), FinVis (Wang et al., 2023b), AI2D (Kembhavi et al., 2016), SciVerse (Guo et al., 2025e), MapQA (Chang et al., 2022) and TQA (Kembhavi et al., 2017). However, FinVis is a
@@
Figure 2: The questioner-inspector data pipeline for generating StockQA instruction tuning dataset.
@@
caption dataset in Chinese, which may result in a language gap and is not convenient for evaluation. Therefore, we regenerate the SFT and test data as multi-choice questions and yes-or-no questions using a questioner-inspector data pipeline. Fig.˜2 shows the overall data pipeline. We use two agents, a QA generator and an inspector. Considering the varying task difficulties, we use Qwen2.5-VL-72b (Bai et al., 2025) to generate multiple choice QA pairs and Qwen2.5-VL-7b to generate Y/N QA pairs. For the inspector, we use Qwen2.5-VL-7b to check the correctness of each QA pair. After initial inspection, rule-based formatting is applied to generate the final dataset, named StockQA. All experiments are conducted using the vllm (Kwon et al., 2023) engine. Appendix˜B provides detailed prompts for each agent, rules for filtering, examples, and statistics of the StockQA dataset. Tab.˜1 shows the statistics of the datasets for DCL and Fig.˜1 shows some examples. More examples are provided in the Sec.˜G.1.
@@
Ability Continual Learning (ACL). DCL assumes that training and test data are IID. However, achieving IID between training and test sets is often challenging in real-world scenarios, which has been ignored by existing benchmarks (Chen et al., 2024a; Zeng et al., 2024; Guo et al., 2025a; Cao et al., 2024). Therefore, we consider a more challenging setting with non-IID training and test data, which we term ability continual learning. For ACL, we select four fundamental abilities for the MLLM to learn sequentially: OCR, math & logic, visual perception, and GUI agent. In terms of the SFT data, we collect the training data from LLaVA-OneVision (Li et al., 2024a), Monkey (Li et al., 2024b), ScreenQA (Hsiao et al., 2022), Screen2Words (Wang et al., 2021), MultiUI (Liu et al., 2024b), Math-LLaVA (Shi et al., 2024b), MAVIS (Zhang et al., 2024b), CLVER (Johnson et al., 2017) and TallyQA (Acharya et al., 2019) and testing data from OCRBench (Liu et al., 2024d), MathVista (Lu et al., 2024), MMTBench-GUI (Ying et al., 2024) and CV-Bench-Counting (Tong et al., 2024), respectively. Tab.˜1 presents the details of the datasets for training and testing in ACL, and Fig.˜1 provides a demonstration. Additional examples can be found in the Sec.˜G.1.
@@ 4 The Proposed Method: MR-LoRA
4.1 Training: Expert Learning without Task Conflict
@@
Figure 4: Comparison of new task performance (LLaVA-based) on both domain and ability CL.
Learning Low-Rank Expert without Task Conflict. In traditional continual learning, particularly class-incremental learning, the model for learning a new task is typically initialized with parameters from the previous task to facilitate knowledge transfer, and then various regularization constraints are incorporated to mitigate catastrophic forgetting. Therefore, a natural question arises: Is this paradigm suitable for continual learning in MLLMs? Some studies (Wei et al., 2025; Yang et al., 2024) have revealed that data interference widely exists in the training of MLLMs. We empirically investigate the task conflict problem of domain and ability continual learning by comparing the average new task performance. The results in Fig.˜4 yield the following observation: (1) Initializing with weights from prior tasks (e.g., LoRA-FT, MoELoRA (Chen et al., 2024a)) reduces model plasticity, leading to worse performance than learning each task individually with randomly initialized LoRA (i.e., scratch). (2) Regularization (e.g., O-LoRA (Wang et al., 2023a), SEFE (Chen et al., 2025)) or parameter-sharing-based methods (e.g., CL-MoE (Huai et al., 2025), HiDE (Guo et al., 2025a)) also suffer from loss of plasticity when learning new tasks. (3) The task conflict in DCL is more severe than that in ACL, which is reasonable because the domain gap in DCL (e.g., autonomous driving vs. science) is often larger than that in ACL (OCR vs. Math). Based on the above analysis, we propose initializing a fresh LoRA (Hu et al., 2021) module from scratch for each task to circumvent inter-task conflicts when learning new domains. Compared to the original parameters of the large model, LoRA introduces minimal additional parameters, enabling domain-specific adaptation via lightweight, task-exclusive adapters.
Few-shot Router Tuning. In our framework, we tune a low-rank expert for each domain or capability, and dynamically select the most appropriate expert at inference time. While existing selection strategies (Zeng et al., 2024; Guo et al., 2025a) rely on simple similarity measures, e.g., computing cosine similarity between task prototypes and sample features in the embedding space, multimodal scenarios involve more complex inputs. Therefore, we propose leveraging the MLLM’s intrinsic capability to process complex multimodal inputs by tuning an MLLM-based selection router. This router identifies the corresponding expert for each input. Specifically, for each task, we collect a few-shot set $\mathcal{M}_{t}=\{(x^{\text{img}}_{t,i},x^{\text{ins}}_{t,i})\}_{i=1}^{m}$, where $m\ll N_{t}$ (we set $m=20$ in all experiments). After each continual learning phase, the accumulated few-shot data $\{\mathcal{M}_{j}\}_{j=1}^{t}$ and expert model descriptions are transformed into structured instructions. We adopt a generative style to select the most suitable expert and tune the MLLM using a router LoRA via autoregressive loss (Liu et al., 2024a). An illustration of the router selection prompt for domain continual learning is provided in Fig.˜3.
4.2 Inference: Router Selection with MLLM